# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ V.4 Multiparty splitting technique
where $\circ$ denotes composition of channels. Throughout, we use the notation $\mathcal{D}^{X}[\rho^{X}]=\pi^{X}\operatorname{Tr}[\rho^{X}]$ for the depolarizing channel on some set of degrees of freedom $X$, with $\pi^{X}=\mathbbm{I}_{X}/d_{X}$ the maximally mixed state. We have also temporarily suppressed the average over circuits $\mathbbm{E}_{U\sim\mathcal{U}}$ for notational convenience.
@@
This splitting into regions $I$ and $I^{c}$ is reminiscent of the Ising spins that we encountered in our calculation of quasientropy in Section IV.2: To each choice of $I$ we can associate variables $\sigma_{i}=\pm 1$ living on each site, where $\sigma_{i}=+1$ if $i\notin I$ corresponds to applying $\hat{\mathcal{D}}_{i}$ and $\sigma_{i}=-1$ for $i\in I$ corresponds to applying $\Theta_{i}$. The sum over $I$ in Eq. (39) then becomes a sum over all spin configurations. We can take this line of reasoning one step further by defining a map $\Theta_{A}$ analogous to (40)
$\Theta_{A}$ $=(\text{id}^{A^{\prime}}-\mathcal{D}^{A^{\prime}})\circ\mathcal{T}^{A\rightarrow A^{\prime}M_{A}}.$ (41)
Since $C$ is traced out and the above map is applied to $A$ in Eq. (38), we can treat $A$ and $C$ as two extra sites in addition to those in $B$, which host Ising spins fixed as $\sigma_{A}=-1$, $\sigma_{C}=+1$, which again reminds us of the calculation of quasientropy. Let $\mathcal{I}$ be the set of spin configurations on $ABC$ with $\sigma_{A,C}$ fixed in this way. We then have
@@
where $X_{I}=\Theta_{I}\otimes\hat{D}_{I^{c}}[\Psi^{ABC}]$ is the summand in (42), to obtain a sum over scalars. Upon bounding each norm separately, one arrives at an upper bound for $\overline{\epsilon}$. Unfortunately, if we take this approach here, the bounds we find become vacuous in the thermodynamic limit, unless we take $q$ to scale with $L_{B}$. As we show in Appendix B, for the holographic model the bound we get on $\overline{\epsilon}$ is at least as large as $\sqrt{d^{\prime}\mathcal{Z}_{-+}}$, where $\mathcal{Z}_{-+}$ is the partition function defined in Eq. (21) (without the denominator $\mathcal{Z}_{++}$). Since the free energy density of the Ising model is finite for any nonzero temperature (here given by the bond entanglement entropy), this partition function must scale exponentially with $N$, leading to a vacuous bound.
@@ V.5 Mapping to self-avoiding walks
In this section, we analyse the expression Eq. 42 in more detail. Instead of applying the triangle inequality as in Eq. 43, we group terms that have the same domain walls and bound their norm in terms of the Renyi-2 entanglement entropy across the domain wall. This allows us to bound $\bar{\epsilon}$ in terms of the partition function of self-avoiding walks on the dual lattice $\Lambda^{*}$. We formalize this in the following Lemma, which we prove in this section.
Lemma 7.
The Boltzmann factor is given by
$H[W]=\frac{1}{2}S^{(2)}_{W},$ (46)
We then combine Lemmas 6 and 7, which gives us a bound on MIE as a function of $d^{\prime}$. Writing $F=-\log\mathcal{Z}_{\rm SAW}$ for the free energy of the statistical mechanics model, the optimal value for $d^{\prime}$ is $d^{\prime}=\lceil F^{-2}e^{2F}\rceil$, which ensures that $\log d^{\prime}\geq 2F-2\log(F)$. If we make the assertion that $F\geq 2$, then we have $\overline{\epsilon}\leq\sqrt{F^{-2}e^{2F}+1}(\exp(-e^{F})-1)\leq 2/F$, which allows us to use the inequality $h_{2}(x)\leq 2x\log(1/x)$ for $x\leq 1/2$. Combining all the above we get the following.
Theorem 8.
If |$\ket{\Psi_{\Lambda}}$ obeys the strictly local property (Definition 4) on a planar graph $\Lambda$ and each site is measured with a POVM that forms a 2-design, with outcome $s_{B}$, then the average post-measurement entanglement entropy can be lower bounded as
$\sum_{s_{B}}p_{s_{B}}S(\rho^{A}_{s_{B}})$ $\geq 2F_{\rm SAW}-2\log(eF_{\rm SAW})$ (47)
provided $F_{\rm SAW}\geq 2$, where $F_{\rm SAW}=-\log\mathcal{Z}_{\rm SAW}$ is the free energy of a statistical mechanics model of self-avoiding walks on the dual lattice which separate the region $A$ from $C$ [defined in Eqs. (45, 46)].
In the rest of this section, we prove Lemma 7. As with the calculation of quasientropy, the Ising spins defined in the previous section are subject to boundary conditions on $A$ and $C$. These ensure that every configuration has at least one domain wall separating $A$ from $C$ with $\sigma_{i}=+1$ on the $C$ side and $\sigma_{i}=-1$ on the $A$ side, and which terminates on the boundaries of the system. We refer to these domain walls as separating domain walls, an example of which is depicted in Fig. 6. The intuition we have developed so far indicates that separating domain walls are associated with entanglement, and thus it makes sense to treat them more explicitly. In the following, we will consider pre-measurement states on the triangular lattice, understanding that the square lattice corresponds to the special case where we set the virtual dimension of certain bonds to 1 (those that are horizontal in Figure 6).
@@
For completeness, we prove the above statement for any 2-design POVM in Appendix C.
Finally, for the map $\tilde{\Theta}_{A}$ we can fix the state to be $\sigma^{M_{A}A^{\prime}}=\pi^{A^{\prime}}\otimes\omega^{M_{A}}$, where $\omega^{M_{A}}=\text{diag}(q_{r_{A}})$ is the distribution of $r_{A}$ when applied to $\pi^{A^{\prime}}$, in which case calculations in Ref. [27] show that
$\|\tilde{\Theta}_{A}[X]\|_{2}$ $\leq\sqrt{d^{\prime}}\|X\|_{2}\;\;\forall X$ (62)
Equations (61, 62) can be thought of as bounds on the induced norm $\|\Phi\|_{2\rightarrow 2}\coloneqq\sup_{X}\|\Phi[X]\|_{2}/\|X\|_{2}$ of the channels $\tilde{\Theta}_{i}$, $\tilde{\Theta}_{A}$. Since the $2\rightarrow 2$ norm is multiplicative under tensor products, we can apply each of these inequalities sequentially for each site $i\in W^{-}$ and on $A$ to Eq. (58), to yield $\|T_{r}\|_{1}\leq\sqrt{d^{\prime}}\|\sigma^{AW^{-}}_{r}\|_{2}$. At this point, we use Eq. 54 to reach
@@
Having bounded the norm of $X(\mathcal{I}_{W})$ [Eq. (49)], where $\mathcal{I}_{W}$ is the subset of all spin configurations that host a particular separating domain wall $W$, we are clearly close to the desired bound on our expression for $\overline{\epsilon}$ in Eq. (42). It is tempting to write $\sum_{\mathcal{I}}(\cdots)=\sum_{W:\text{SDW}}\sum_{\mathcal{I}_{W}}(\cdots)$, where the sum over $W$ is over all separating domain walls. By the triangle inequality, this would yield a bound $\overline{\epsilon}\leq\sum_{W:\text{SDW}}\|X(\mathcal{I}_{W})\|_{1}$, where the sum is over all separating domain walls $W$. However, this actually double counts spin configurations that contain more than one separating domain wall, and because the operators $X(\mathcal{I}_{W})$ are not positive semi-definite this means such a bound is not necessarily valid. This issue is resolved with in Appendix D. As a result, we obtain Eq. 44, which differs from the naïve expression onlt by a multiplicatibe factor $f(\mathcal{Z}_{\rm SAW})$, which in the regime of interest $10<\mathcal{Z}_{\rm SAW}<$ 1 is mild, namely $21<f(\mathcal{Z}_{\rm SAW})<$ 2.
Thus we a finally left with a sum over all separating domain walls. By definition, these are all non-intersecting, so we can interpret the set of all $W$ as a collection of self-avoiding walks on the dual lattice $\Lambda^{*}$, leading to Lemma 7. Self-avoiding walks is a well-studied statistical mechanics model in the field of polymer physics, and there exist rigorous results that we use later to obtain a bound in specific cases.
If the original lattice $\Lambda$ is a planar graph that is not a triangulation, then we have the added complication that there can be faces where domain walls intersect. As explained previously, we can always add edges to $\Lambda$ to make a triangulation $\Lambda^{\prime}$, while ensuring that the strictly local condition Definition 4 still holds with respect to $\Lambda^{\prime}$. This will yield a bound on $\overline{\epsilon}$ in terms of self-avoiding walks on the dual lattice of $(\Lambda^{\prime})^{*}$, which may be distinct from $\Lambda^{*}$. For each self-avoiding walk on $(\Lambda^{\prime})^{*}$, there is a corresponding self-avoiding walk on $\Lambda^{*}$, and because the Boltzmann weights only depend on the Rényi entanglement entropy across $W$ (which does not change when extra fictitious bonds are added), we can substitute the partition function for the original dual lattice $\Lambda^{*}$.
@@ VI.1 Random holographic tensor network states
For random holographic circuits with bond state |$\ket{\omega_{e}}$ on a 2D lattice $\Lambda$, with $L_{y}=\textup{poly}(L_{x})$, we can find a constant $\kappa$ such that the average measurement-induced entanglement between $A$ and $C$ in the geometry shown in Fig. 5 satisfies
@@
for $L_{x}$ exceeding some constant. We can take $\kappa$ arbitrarily close to
$\kappa\rightarrow S_{e}^{(2)}-2\log\mu_{\Lambda^{*}}$ (66)
@@
This result establishes that if the entanglement of the pre-measurement state is above some threshold, then extensive long-ranged MIE is produced. For example, for the square lattice, there are known upper bounds on the connective constant $\log\mu_{\Lambda^{*}}\leq 0.97$ [71], yielding an upper bound on the threshold value of the Rényi-2 bond entropy $S_{e,{\rm crit}}\leq 1.94\text{ nats}=2.80\ldots\text{ bits}$. While this should not be thought of as an exact calculation of the critical bond entanglement $S_{e,{\rm crit}}$, due to the fact we have derived only a one-sided bound, we can compare this to the numerically observed value of just over 1 bit [72]. More broadly, this gives a concrete proof that a fixed (system-size independent) amount of entanglement on each bond suffices to generate highly entangled holographic states. In contrast, all previous rigorous estimates assumed a bond dimension growing in some way with system size.
Provided one has the means to evaluate the appropriate generating function $\mathcal{Z}_{\rm SAW}[\beta]$, the above can be quickly generalized to different geometries of the regions $A$, $B$, $C$: The domain wall configurations included are always those that separate $A$ from $C$.
@@ VI.2 Random 4-local square lattice circuits
The proof of this statement is given in Appendix E. In brief, we can use Eq. (55) as a starting point, since no assumptions about the POVMs had been made at that point (only that the state obeyed the strict locality condition), and look in more detail at the intra-cell structure of the measurement maps $\Theta_{i}$. We arrive at an analogous mapping to a self-avoiding walk partition function $\mathcal{Z}_{\rm SAW}$, with slightly modified parameters, and so we can use the same rigorous statistical mechanics results as before to get the bound (67). The specific inequalities we have employed to characterise the intra-cell structure of $\Theta_{i}$ are fairly crude, which is why we end up with a quantitatively conservative restriction on $q$ in the above. While we should not expect the actual critical Hilbert space dimension to be close to 88, importantly we can infer the existence of extensive long-ranged post-measurement entanglement in circuits with constant Hilbert space dimension and constant depth, which is a significant improvement over previous bounds which required $q$ to grow with system size.
As before, the transition of the self-avoiding walk partition function is a bulk property, and thus we expect to see similar kinds of long-ranged MIE for different geometries $A$, $B$, $C$, provided $A$ and $C$ make up a vanishing fraction of the system.
@@ VII.1 Monitored quantum dynamics from sampling shallow circuits and spacetime duality
Figure 8: (a) Starting from an isoTNS, we measure degrees of freedom in $B$. (b) Applying a measurement to a physical degree of freedom converts the local isometric tensor to an unravelled quantum channel, described by a properly normalized Kraus operators $K_{s}$ which act on virtual degrees of freedom. (c) Concatenating the virtual legs together, we obtain a (1+1)D monitored quantum dynamics with a brickwork architecture, with the time-like dimension (here $y$) determined by the direction of the isometries. (d) In the distillation protocol, a projection $L_{r}^{\dagger}$ is applied to $A$ to get a state on a smaller Hilbert space $A^{\prime}$. Within this subspace, the sideways-evolving dynamics can be approximated by an ensemble of effective unitaries $U^{\rm eff}_{s_{B}}$ which depend on the random outcomes $s_{B}$.
@@
Figure 8(a) illustrates the setup we consider, where one performs projective measurements on an isoTNS with the regions $ABC$ chosen as in Fig. 1. The significance of the isometric property of the local tensors is that isometries represent physical quantum operations. In this case, each tensor constitutes a map from two virtual legs $i_{1}i_{2}$ to the physical leg $Q$ plus the remaining two virtual legs $o_{1}o_{2}$, and in principle this map could be implemented using unitaries and ancillas. If we append a measurement to the physical leg $Q$, which itself is also a physical operation, then we obtain a unravelled quantum channel from $i_{1}i_{2}$ to $o_{1}o_{2}$ (see Ref. [82]). These unravelled channels, which are specified by a set of Kraus operators $\{K_{s}\}$, describe stochastic quantum processes where the state $\rho$ evolves to $\rho_{s}^{\prime}=K_{s}\rho K_{s}^{\dagger}/p(s|\rho)$ with probability $p(s|\rho)=\operatorname{Tr}[K_{s}^{\dagger}K_{s}\rho]$. The Kraus operators for a given site $i$ can be related to the local tensor by projecting out the physical leg $K_{s_{i}}=(\bra{s_{i}}_{Q_{i}}\otimes\mathbbm{I}_{o_{1}o_{2}})V^{i_{1}i_{2}\rightarrow Po_{1}o_{2}}$ [Fig. 8(b)]. Using the fact that $V^{i_{1}i_{2}\rightarrow Po_{1}o_{2}}$ is an isometry, one can verify that the unravelled map

### v2

@@ V.4 Multiparty splitting technique
where $\circ$ denotes composition of channels. Throughout, we use the notation $\mathcal{D}^{X}[\rho^{X}]=\pi^{X}\operatorname{Tr}[\rho^{X}]$ for the depolarizing channel on some set of degrees of freedom $X$, with $\pi^{X}=I_{X}/d_{X}$ the maximally mixed state. We have also temporarily suppressed the average over circuits $\mathbbm{E}_{U\sim\mathcal{U}}$ for notational convenience.
@@
This splitting into regions $I$ and $I^{c}$ is reminiscent of the Ising spins that we encountered in our calculation of quasientropy in Section IV.2: To each choice of $I$ we can associate variables $\sigma_{i}=\pm 1$ living on each site, where $\sigma_{i}=+1$ if $i\notin I$, and $\sigma_{i}=-1$ if $i\in I$; these variables signify which of the maps $\hat{\mathcal{D}}_{i}$ and $\Theta_{i}$ is applied to site $i$. The sum over $I$ in Eq. (39) then becomes a sum over all spin configurations.
We can build on this interpretation by observing that the overall map applied to region $A$ in Eq. (38) bears a resemblance to $\Theta_{i}$. Indeed, $\mathcal{T}^{Q_{i}\rightarrow M_{i}}$ can be thought of as an instance of the distillation channel (2) in the special case of fully destructive measurements $d^{\prime}$. It is therefore natural to define the analogous map
$\Theta_{A}$ $\coloneqq(\text{id}^{A^{\prime}}-\mathcal{D}^{A^{\prime}})\circ\mathcal{T}^{A\rightarrow A^{\prime}M_{A}}.$ (41)
Noticing also that region $C$ is traced out in (38), it is helpful to view $A$ and $C$ as two extra sites in addition to those in $B$, which host Ising spins fixed as $\sigma_{A}=-1$, $\sigma_{C}=+1$. This again reminds us of the calculation of quasientropy. Let $\mathcal{I}$ be the set of spin configurations on $ABC$ with $\sigma_{A,C}$ fixed in this way. We then have
@@
where $X_{I}=\Theta_{I}\otimes\hat{D}_{I^{c}}[\Psi^{ABC}]$ is the summand in (42), to obtain a sum over scalars. Upon bounding each norm separately, one arrives at an upper bound for $\overline{\epsilon}$. Unfortunately, if we take this approach here, the bounds we find become vacuous in the thermodynamic limit, unless we take $q$ to scale with $L_{B}$. As we show in Appendix B, for the holographic model the bound we get on $\overline{\epsilon}$ is at least as large as $\sqrt{d^{\prime}\mathcal{Z}_{-+}}$, where $\mathcal{Z}_{-+}$ is the partition function defined in Eq. (21) (without the denominator $\mathcal{Z}_{++}$). Since the free energy density of the Ising model is finite for any nonzero temperature (here given by the bond entanglement entropy), this partition function must scale exponentially with $N$, leading to a vacuous bound. Thus, we will have to take a different approach.
@@ V.5 Mapping to self-avoiding walks
Instead of applying the triangle inequality as in Eq. 43, in this section we analyse the expression Eq. 42 more carefully. This will result in a nontrivial upper bound on $\epsilon$ in terms of a different statistical mechanics partition function, which describes self-avoiding walks, Lemma 7. In turn, we will arrive at our main technical result, Theorem 8, a lower bound on MIE.
The Ising spins defined in the previous section are reminiscent of those that appeared in the calculation of quasientropy. As before, they are subject to boundary conditions on $A$ and $C$. These ensure that every configuration has at least one domain wall separating $A$ from $C$ with $\sigma_{i}=+1$ on the $C$ side and $\sigma_{i}=-1$ on the $A$ side, and which terminates on the boundaries of the system. We refer to these domain walls as separating domain walls, an example of which is depicted in Fig. 6. The intuition we have developed so far indicates that separating domain walls are associated with entanglement, and thus it makes sense to treat them more explicitly. In the following, we will consider pre-measurement states on the triangular lattice, understanding that the square lattice corresponds to the special case where we set the virtual dimension of certain bonds to 1 (those that are horizontal in Figure 6).
@@
We note that the proof of Eq. 57, which is given in Appendix C for completeness, is the only point in our argument that requires the circuit to be random.
Finally, for the map $\tilde{\Theta}_{A}$ we can fix the state to be $\sigma^{M_{A}A^{\prime}}=\pi^{A^{\prime}}\otimes\omega^{M_{A}}$, where $\omega^{M_{A}}=\text{diag}(q_{r_{A}})$ is the classical distribution of unitaries $V_{r_{A}}$ appearing in the definition of the distillation channel (2). Because these unitaries $V_{r_{A}}$ form a 2-design we can use the approach given in Ref. [27] (see Appendix C) to obtain an analogous result
$\|\tilde{\Theta}_{A}[X]\|_{2}$ $\leq\sqrt{d^{\prime}}\|X\|_{2}\;\;\forall X.$ (58)
These inequalities can be viewed as bounds on the induced norm $\|\Phi\|_{2\rightarrow 2}\coloneqq\sup_{X}\|\Phi[X]\|_{2}/\|X\|_{2}$ of the channels $\tilde{\Theta}_{i}$, $\tilde{\Theta}_{A}$. Since the $2\rightarrow 2$ norm is multiplicative under tensor products, we can apply each of these inequalities sequentially for each site $i\in W^{-}$ and on $A$ to Eq. (54), to yield $\|T_{r}\|_{1}\leq\sqrt{d^{\prime}}\|\sigma^{AW^{-}}_{r}\|_{2}$. At this point, we use Eq. 50 to reach
@@
Having bounded the norm of $X(\mathcal{I}_{W})$ [Eq. (45)], where $\mathcal{I}_{W}$ is the subset of all spin configurations that host a particular separating domain wall $W$, we are clearly close to the desired bound on our expression for $\overline{\epsilon}$ in Eq. (42). It is tempting to write $\sum_{\mathcal{I}}(\cdots)=\sum_{W:\text{SDW}}\sum_{\mathcal{I}_{W}}(\cdots)$, where the sum over $W$ is over all separating domain walls. By the triangle inequality, this would yield a bound $\overline{\epsilon}\leq\sum_{W:\text{SDW}}\|X(\mathcal{I}_{W})\|_{1}$, where the sum is over all separating domain walls $W$. Since domain walls take the form of self-avoiding walks on the dual lattice $\Lambda^{*}$, this would then yield the desired bound on $\bar{\epsilon}$ in terms of a self-avoiding walk partition function.
In fact, this naïve expression double counts spin configurations that contain more than one separating domain wall. Because the operators $X(\mathcal{I}_{W})$ are not positive semi-definite this means such a bound is not necessarily valid. This issue is resolved in Appendix D, where we prove that such configurations only lead to a subleading correction. As a result, we obtain the following.
Lemma 7.
The Boltzmann factor is given by
$H[W]=\frac{1}{2}S^{(2)}_{W},$ (62)
Eq. 60 includes a multiplicative factor $f(\mathcal{Z}_{\rm SAW})$ coming from configurations with multiple separating domain walls. In the regime of interest $10<\mathcal{Z}_{\rm SAW}<$ 1, this correction is mild, namely $21<f(\mathcal{Z}_{\rm SAW})<$ 2. Self-avoiding walks is a well-studied statistical mechanics model in the field of polymer physics, and there exist rigorous results that we use later to obtain a bound in specific cases.
Finally, we can combine Lemmas 6 and 7, which gives us a bound on MIE for any arbitrary integer choice of $d^{\prime}$, namely
$\overline{E(A:C)}$ $\geq\left(1-\frac{\sqrt{d^{\prime}}g(F)}{2}\right)\log d^{\prime}-h_{2}\left(\frac{\sqrt{d^{\prime}}g(F)}{2}\right)$ (63)
where $F=-\log\mathcal{Z}_{\rm SAW}$ is the free energy of the statistical mechanics model defined above, and $g(x)\coloneqq\exp(e^{-x})-1$. We now look to optimize our choice of $d^{\prime}$ to obtain the tightest possible bound. A simple yet near-optimal choice is $d^{\prime}=\lceil F^{-2}e^{2F}\rceil$, which is exponentially large in $F$ while still ensuring that $\bar{\epsilon}$ decays with system size. If we make a further assertion that $F\geq 2$, then $\sqrt{d^{\prime}}g(F)\leq\sqrt{F^{-2}e^{2F}+1}(1-\exp(-e^{F}))\leq 2/F$, and we can then use the inequality $h_{2}(x)\leq 2x\log(1/x)$ for $x\leq 1/2$. Combining all the above we have
Theorem 8.
If |$\ket{\Psi_{\Lambda}}$ obeys the strictly local property (Definition 4) on a planar graph $\Lambda$ and each site in region $B$ is measured with a POVM that forms a 2-design, with outcome $s_{B}$, then the average post-measurement entanglement entropy can be lower bounded as
$\sum_{s_{B}}p_{s_{B}}S(\rho^{A}_{s_{B}})$ $\geq 2F_{\rm SAW}-2\log(eF_{\rm SAW})$ (64)
provided $F_{\rm SAW}\geq 2$, where $F_{\rm SAW}=-\log\mathcal{Z}_{\rm SAW}$ is the free energy of a statistical mechanics model of self-avoiding walks on the dual lattice which separate the region $A$ from $C$ [defined in Eqs. (61, 62)].
@@ VI.1 Random holographic tensor network states
For random holographic circuits with bond state |$\ket{\omega_{e}}$ on a 2D lattice $\Lambda$ of dimensions $L_{x}\times L_{y}$, with $L_{y}=\textup{poly}(L_{x})$, the average measurement-induced entanglement between $A$ and $C$ in the geometry shown in Fig. 5 satisfies
@@
for $L_{x}$ exceeding some constant, and a constant $\kappa$ that can be taken to be arbitrarily close to
$\kappa\rightarrow S_{e}^{(2)}-2\log\mu_{\Lambda^{*}},$ (67)
@@
This result establishes that if the entanglement of the pre-measurement state is above some threshold, then extensive long-ranged MIE is produced. For example, for the square lattice, there are known upper bounds on the connective constant $\log\mu_{\Lambda^{*}}\leq 0.97$ [71], yielding an upper bound on the threshold value of the Rényi-2 bond entropy
$S_{e,{\rm crit}}$ $\leq 2.80\text{ bits}.$ (square lattice) (68)
The corresponding values for different lattices can be obtained from estimates of the corresponding connective constants. More broadly, this gives a concrete proof that a fixed (system-size independent) amount of entanglement on each bond suffices to generate highly entangled holographic states. In contrast, all previous rigorous estimates assumed a bond dimension growing in some way with system size.
Provided one has the means to evaluate the appropriate generating function $\mathcal{Z}_{\rm SAW}[\beta]$, the above can be quickly generalized to different geometries of the regions $A$, $B$, $C$: The domain wall configurations included are always those that separate $A$ from $C$. Given that the transition of the self-avoiding walk partition function is a bulk property, we expect that long-ranged MIE will arise in the regime (68) for different geometries $A$, $B$, $C$, provided $A$ and $C$ make up a vanishing fraction of the system.
@@
Eq. 68 gives us a sufficient condition for long-ranged MIE to arise in random holographic circuits, and we can ask how closely this condition describes the true critical point. Numerical studies of random tensor networks in a square lattice geometry (which are equivalent to holographic circuits with maximally entangled bond states; see Section III.3) indicate that the critical bond dimension is a little above 2, i.e. $S_{e,{\rm crit}}\gtrsim 1\text{ bit}$ [72]. Thus, our bounds give an estimate of $S_{e,{\rm crit}}$ of the correct order of magnitude.
@@ VI.2 Random 4-local square lattice circuits
The proof of this statement is given in Appendix E. In brief, we can use Eq. (51) as a starting point, since no assumptions about the POVMs had been made at that point (only that the state obeyed the strict locality condition), and look in more detail at the intra-cell structure of the measurement maps $\Theta_{i}$. We arrive at an analogous mapping to a self-avoiding walk partition function $\mathcal{Z}_{\rm SAW}$, with slightly modified parameters, and so we can use the same rigorous statistical mechanics results as before to get the bound (69).
The specific inequalities we have employed to characterise the intra-cell structure of $\Theta_{i}$ are fairly crude. This is why we obtain a relatively conservative bound on $q$, whereas we would expect the true critical value to be somewhat similar to the numerically observed value for the holographic architecture $q_{\rm crit}\gtrsim 2$ [72]. Even so, the key insight from Proposition 10 is that we can infer the existence of extensive long-ranged post-measurement entanglement in circuits with constant Hilbert space dimension and constant depth, which is a significant improvement over previous bounds which required $q$ to grow with system size.
@@ VII.1 Monitored quantum dynamics from sampling shallow circuits and spacetime duality
Figure 8: (a) Starting from an isoTNS, we measure degrees of freedom in $B$. (b) Applying a measurement to a physical degree of freedom converts the local isometric tensor to an unravelled quantum channel, described by properly normalized Kraus operators $K_{s}$ which act on virtual degrees of freedom. (c) Concatenating the virtual legs together, we obtain a (1+1)D monitored quantum dynamics with a brickwork architecture, with the time-like dimension (here $y$) determined by the direction of the isometries. (d) In the distillation protocol, a projection $L_{r}^{\dagger}$ is applied to $A$ to get a state on a smaller Hilbert space $A^{\prime}$. Within this subspace, the sideways-evolving dynamics can be approximated by an ensemble of effective unitaries $U^{\rm eff}_{s_{B}}$ which depend on the random outcomes $s_{B}$.
@@
Figure 8(a) illustrates the setup we consider, where one performs projective measurements on an isoTNS with the regions $ABC$ chosen as in Fig. 1. The significance of the isometric property of the local tensors is that isometries represent physical quantum operations. In this case, each tensor constitutes a map from two virtual legs $i_{1}i_{2}$ to the physical leg $Q$ plus the remaining two virtual legs $o_{1}o_{2}$, and in principle this map could be implemented using unitaries and ancillas. If we append a measurement to the physical leg $Q$, which itself is also a physical operation, then we obtain a unravelled quantum channel from $i_{1}i_{2}$ to $o_{1}o_{2}$ (see Ref. [82]). These unravelled channels, which are specified by a set of Kraus operators $\{K_{s}\}$, describe stochastic quantum processes where the state $\rho$ evolves to $\rho_{s}^{\prime}=K_{s}\rho K_{s}^{\dagger}/p(s|\rho)$ with probability $p(s|\rho)=\operatorname{Tr}[K_{s}^{\dagger}K_{s}\rho]$. The Kraus operators for a given site $i$ can be related to the local tensor by projecting out the physical leg $K_{s_{i}}=(\bra{s_{i}}_{Q_{i}}\otimes I_{o_{1}o_{2}})V^{i_{1}i_{2}\rightarrow Po_{1}o_{2}}$ [Fig. 8(b)]. Using the fact that $V^{i_{1}i_{2}\rightarrow Po_{1}o_{2}}$ is an isometry, one can verify that the unravelled map